OpenStudy (anonymous):

CAN ANYONE help me with this equation? turn from ration to a quadratic equation then solve: a.) 3/2-x + 3/2+x=4 b.) 1/x+3 + 1/x = 1/2

OpenStudy (anonymous):

\[\frac{ 3 }{ 2-x }+\frac{ 3 }{ 2+x }=4\] \[3\left[ \frac{ 1 }{ 2-x }+\frac{ 1 }{ 2+x } \right]=4\] \[3\left[ \frac{ 2+x+2-x }{ \left( 2-x \right)\left( 2+x \right) } \right]=4\] \[\frac{ 3*4 }{ 4 }=4-x^2,x^2=4-3=1,x=\pm 1\]

OpenStudy (anonymous):

thank you thank you

OpenStudy (anonymous):

yw